I hope this is the right section! Um...here is my question:
x-2/x + 4/5x = -1/5 ... the / are like saying x-2 over x
This is what I tried:
5x(x-2) + 4(x)/5x^2 = -1/5
-6x = -1/5
x = 1/30
I don't know what I did wrong because my textbook said the answer was 1. If I could have help that would be great! Thanks!

Answer by Alan3354(69443)

x-2/x + 4/5x = -1/5 ... the / are like saying x-2 over x
This is what I tried:
5x(x-2) + 4(x)/5x^2 = -1/5
-6x = -1/5
x = 1/30
------------
(x-2)/x + 4/(5x) = -1/5 (I think this is what you meant)
Multiply thru by 5x to clear fractions
5(x-2) + 4 = -x
5x-10 + 4 = -x
6x = 6
x = 1